# Seungone Kim1,2,3  Juyoung $\textbf{}^{*}$  Shayne Longpre4Bill Yuchen Lin5Jamin Shin1 Sean Welleck3 Graham Neubig3Moontae Lee2,6Kyungjae Lee2Minjoon Seo1

### v1

@@ Abstract
Proprietary LMs such as GPT-4 are often employed to assess the quality of responses from various LMs. However, concerns including transparency, controllability, and affordability strongly motivate the development of open-source LMs specialized in evaluations. On the other hand, existing open evaluator LMs exhibit critical shortcomings: 1) they issue scores that significantly diverge from those assigned by humans, and 2) they lack the flexibility to perform both direct assessment and pairwise ranking, the two most prevalent forms of assessment. Additionally, they do not possess the ability to evaluate based on custom evaluation criteria, focusing instead on general attributes like helpfulness and harmlessness. To address these issues, we introduce Prometheus 2, a more powerful evaluator LM than it’s predecessor that closely mirrors human and GPT-4 judgements. Moreover, it is capable of processing both direct assessment and pair-wise ranking formats grouped with a user-defined evaluation criteria. On four direct assessment benchmarks and four pairwise ranking benchmarks, Prometheus 2 scores the highest correlation and agreement with humans and proprietary LM judges among all tested open evaluator LMs. Our models, code, and data are all publicly available  [^1] https://github.com/prometheus-eval/prometheus-eval.
@@ 1 Introduction
However, relying on proprietary LMs for evaluation poses significant challenges. The lack of transparency about their training data compromises both fairness and compliance, making it problematic to use them in evaluation pipelines. Additionally, concerns regarding controllability and affordability also persist (Kim et al., 2023). To address these issues, recent works have focused on developing evaluator LMs that are open-access, transparent, and controllable (Kim et al., 2023; Wang et al., 2023a, b; Li et al., 2023a; Zhu et al., 2023; Jiang et al., 2023b, c; Lee et al., 2024). Yet, these models often yield scoring decisions that do not correlate well enough with human judgments or those made by proprietary LMs, failing to effectively simulate them. Moreover, open evaluator LMs are not flexible since they are typically trained only to perform either direct assessment or pairwise ranking and assess based on general public preferences like helpfulness and harmlessness, limiting their ability to handle diverse real-life scenarios.
@@ 2.1 Language Model-based Evaluation
To assess the generation capabilities of LMs, prior works such as the GEM benchmark (Gehrmann et al., 2021, 2022) employed Rouge (Lin, 2004), BLEU (Papineni et al., 2002), and BERTScore (Zhang et al., 2019) as their metric, which measures the lexical or semantic similarity between a reference answer and a response. However, these conventional metrics are prone to false negatives because they are not expressive enough to recognize responses that are of good quality but differ from the reference answer (Schluter, 2017; Freitag et al., 2020; Hanna and Bojar, 2021).
Recently, employing language models as a judge has gained attention as a promising paradigm to mimic the depth and granularity that human evaluation offers (Zheng et al., 2023; Liu et al., 2023b; Li et al., 2023b; Chan et al., 2023; Ye et al., 2023). To reduce the over-reliance on proprietary LMs, follow-up works suggest training language models specialized in evaluations (Cui et al., 2023; Kim et al., 2023; Jiang et al., 2023b, c; Li et al., 2023a; Lee et al., 2024). Yet, open evaluator LMs do not possess the flexibility to function in different evaluation schemes and show weak evaluation performances compared to proprietary LMs. We aim to bridge this gap by introducing Prometheus 2.
@@ 2.2 Weight Merging
Prior works have demonstrated that weight merging can enhance performances across various domains, including language modeling (Li et al., 2022; Matena and Raffel, 2022; Ilharco et al., 2022; Don-Yehiya et al., 2022; Gururangan et al., 2023; Yadav et al., 2024; Sukhbaatar et al., 2024), instruction-tuning (Jang et al., 2023b; Yu et al., 2023), and aligning to user perferences (Jang et al., 2023a; Rame et al., 2024; Wang et al., 2024). In our work, we specifically focus on enhancing the evaluation capabilities of open evaluator LMs. By merging models trained on different assessment formats—specifically, direct assessment and pairwise ranking—we aim to obtain an evaluator LM that not only functions in both formats but also shows as good evaluation performances as proprietary LMs.
@@ 3 Methodology
We propose a new recipe for training a unified evaluator LM based on merging the weights of models trained for direct assessment and pairwise ranking. We begin with background on direct assessment and pairwise ranking for evaluator LMs (Section 3.1, 3.2), followed by the construction process of our training data (Section 3.3). Finally, we present our methods to train the state-of-the-art evaluator LM, Prometheus 2 models (Section 3.4).
@@ 3.1 Direct Assessment
Direct assessment is mapping an instruction $i$ and response $r$ into a scalar value score $s$, such as ${f}_{direct}:(i,r)\mapsto s\text{ where }s\in\mathbb{R}$. For the scoring range, we use a 1-5 Likert scale scoring.
Prior works have identified several recipes to align the scores provided by evaluator LMs (${s}_{LM}$) and the scores assigned by humans (${s}_{human}$). For instance, Liu et al. (2023a) and Zheng et al. (2023) have shown that it is crucial to add a reference answer $a$ as input to the evaluator LM to maximize the correlation between ${s}_{LM}$ and ${s}_{human}$. Also, Zheng et al. (2023) and Ye et al. (2023) showed that prompting the language model to write verbal feedback ${v}_{r}$ before $s$ also improves the correlation between ${s}_{LM}$ and ${s}_{human}$. Lastly, Ye et al. (2023) and Kim et al. (2023) showed that by explicitly integrating evaluation criteria $e$, users can define the standards for model assessment, ensuring evaluations are flexible to specific needs rather than generic qualities. Specifically, $e$ is represented as a score rubric including a description for the criteria itself and a set of descriptions for each score between the scoring range. This is expressed as:
@@ 3.2 Pairwise Ranking
In pairwise ranking, the evaluation criteria $e$ do not include a set of descriptions for each score; instead, only the description of the evaluation criterion itself. Also, it is noteworthy that the verbal feedback ${v}_{{r}_{m},{r}_{n}}$ compares the commonalities and differences between ${r}_{m}$ and ${r}_{n}$ concerning $e$.
@@ 3.3 The Preference Collection
Popular pairwise ranking datasets such as HH-RLHF (Bai et al., 2022) or Ultra Feedback (Cui et al., 2023) do not include an evaluation criteria $e$ and a verbal feedback ${v}_{{r}_{m},{r}_{n}}$. To obtain an evaluator LM that could assess based on what users care about, we construct the Preference Collection that includes 1K evaluation criteria.
Construction Process
To construct the Preference Collection, we apply two modifications to the Feedback Collection. First, since the Feedback Collection includes five responses for each instruction, each corresponding to a scoring decision between 1 and 5, we pair two out of the five responses, resulting in a total of ten combinations per instruction. Using the existing scoring decisions for each response, we determine which response is better and assign a new scoring decision for that pair (i.e., “Response A is better” or “Response B is better”). Second, to generate new verbal feedback ${v}_{{r}_{m},{r}_{n}}$ for each pair of responses, we prompt GPT-4-1106 to identify the commonalities and differences of the two responses.
The statistics of the resulting dataset are listed in Table 1 along with the Feedback Collection. We explain about our quality verification process of the Preference Collection in Appendix A. Also, we include the prompts we use for the augmentation process in Appendix F.
Direct Assessment Benchmarks Pairwise Ranking Benchmarks
Vicuna MT FLASK Feedback HHH MT Bench Auto-J Preference
Bench Bench Bench Align. Human Judg. Eval Bench
Judgment Source Proprietary LMs Proprietary LMs Proprietary LMs Proprietary LMs Humans Humans Humans Proprietary LMs
& Humans
Metrics Correlation Correlation Correlation Correlation Accuracy Accuracy Accuracy Accuracy
Reference Answer Y Y Y Y N N N Y
# Score Rubrics 80 80 12 200 4 1 1 200
# Instructions 80 80 200 200 221 80 58 200
# Judgments 320 320 2,000 1,000 221 3,360 1,392 2,000
@@
3.4 Employing Evaluator Language Models
@@
Prompting involves querying an LM to make judgments in a specified evaluation format without training on any feedback dataset.
@@
Single-Format training involves training a base model $\theta$ on either on a direct assessment feedback dataset ${D}_{d}$ or a pairwise ranking feedback dataset ${D}_{p}$.
@@
Joint training involves training a base model $\theta$ on both a direct assessment feedback dataset ${D}_{d}$ and a pairwise ranking feedback dataset ${D}_{p}$. This enables the resulting evaluator LM to function across both evaluation formats.
@@
Weight Merging involves training two models, ${\theta}_{d}$ and ${\theta}_{p}$, separately on a direct assessment feedback dataset ${D}_{d}$ and a pairwise ranking feedback dataset ${D}_{p}$. Then, we obtain the final evaluator LM ${\theta}_{final}$ with linear merging :
@@
We conduct experiments by using $\alpha=0.5$. In Section 6.3, we observe how altering the coefficient $\alpha$ affects downstream performance on each evaluation scheme. We empirically find that this simple recipe work best when we choose Mistral-7B as our base model. In addition to linear merging, we also test different merging techniques including:
    $(1-\alpha)\times({\theta}_{p}-{\theta}_{init})$
  • DARE merging (Yu et al., 2023), while also similar to Task Arithmetic and TIES merging, performs a Random Drop and Re-scale operations in the task vector ${\theta}_{d}-{\theta}_{init}$ and ${\theta}_{p}-{\theta}_{init}$ to remove redundant weights. We find that DARE merging work best when we choose Mixtral-8x7B as our base model.
@@ 4 Experimental Setup
In this section, we explain our experimental setup to assess evaluator LMs. We first explain the benchmarks and metrics we employ (Section 4.1) and the baselines we use as evaluator LMs (Section 4.2).
4.1 Benchmarks and Metrics
The statistics of all the benchmarks are in Table 2.
The four direct assessment benchmarks are:
@@ 4.1 Benchmarks and Metrics
In direct assessment, we conduct reference-based evaluations by appending the reference answer as the input. We use Pearson, Spearman, and Kendall-Tau as performance metrics to measure scoring correlations against reference evaluators.
@@
4.2 Baselines
Prompting Baselines
We employ Llama-2-Chat-7,13,70B (Touvron et al., 2023); Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a); and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024) as our baselines. It’s worth noting that models not explicitly trained on feedback data often fail to generate responses in the required format, making it extremely difficult to parse scoring decisions. Although it is impractical for regular use, we make a fair comparison by infinitely looping until scores can be parsed. Also, we include proprietary LMs such as GPT-3.5-Turbo-0613; GPT-4-1106; and Claude-3-Opus.
Single-Format Trained Evaluator LMs
For single-format trained evaluator LMs, we test Prometheus-7,13B (Kim et al., 2023) (direct assessment); UltraRM-13B (Cui et al., 2023) (pairwise ranking); and PairRM-0.4B (Jiang et al., 2023c) (pairwise ranking). In addition, we also report the performances of single-format training Mistral-7B-Instruct-v0.2 and Mixtral-8x7B-Instruct-v0.1 on either direct assessment or pairwise ranking.
Jointly Trained Evaluator LMs
For jointly trained evaluator LMs, we test Auto-J (Li et al., 2023a). In addition, we report the performances of jointly training Mistral-7B and Mixtral-8x7B on both direct assessment and pairwise ranking.
Weight Merging
Prometheus 2 (7B & 8x7B) models are our weight merging baselines.
Details on the hyper-parameters for training and inference along with the prompt templates are all listed in Appendix B, G, H.
@@ 4.2 Baselines
Evaluator LM HHH Alignment MT Bench Human Judg. Auto-J Eval
Auto-J (13B) 46.61 75.57 28.96 48.14 69.12 20.98 47.40 76.64 29.24
Prometheus-2-7B 74.21 78.73 4.52 63.24 67.25 4.01 68.11 73.80 5.69
Claude-3-Opus 84.62 94.57 9.95 62.65 77.65 15.00 61.04 82.90 21.86
@@ 5.1 Direct Assessment Results
The direct assessment results are shown in Table 3. The scoring decisions of Prometheus-2 models (7B & 8x7B), GPT-4-1106, Claude-3-Opus, and human evaluators all strongly correlate with each other, yielding Pearson correlations higher than 0.5 regardless of the reference evaluator and benchmark. On the other hand, base LMs, single-format trained LMs, and jointly trained LMs show lower correlations with GPT-4-1106, Claude-3-Opus, and humans, mostly falling below 0.5.
Notably, Prometheus 2 models outperform Prometheus and Auto-J by at least 0.2 units across benchmarks in their correlation with proprietary LMs. Moreover, on the FLASK benchmark, while the correlation between humans and GPT-4 is 0.679, the highest correlation previously achieved by Prometheus-13B with humans was 0.449, but Prometheus-2-8x7B achieves a correlation of 0.555 with humans, effectively halving the gap.
@@ 5.2 Pairwise Ranking Results
The pairwise ranking results are shown in Table 4. We exclude the results of Pair RM, Ultra RM on ‘w/ Tie’ settings since they could not give tie options.
@@
5.3 Consistency Across Evaluation Formats
In addition to obtaining high correlation and accuracy, achieving high consistency is another important aspect for evaluator LMs. Specifically, we conduct an experiment testing if evaluator LMs could achieve consistent scores across different evaluation formats. To do this, we use pairwise ranking benchmarks and measure the performance differences when prompted with direct assessment formats and pairwise ranking formats. Specifically, following Kim et al. (2023), to process pairwise ranking datasets in a direct assessment scheme, we evaluate each response separately and compare the scoring decisions. We mark it as correct if the evaluator LM provides a higher score for the human-chosen response over the rejected one. As shown in Table 5, the results show that Prometheus 2 models show lower performance differences across evaluation formats, indicating their robustness.
Model Direct Assessment Benchmarks Pairwise Ranking Benchmarks
@@ 5.3 Consistency Across Evaluation Formats
Direct Assessment & Pairwise Ranking 0.642 0.543 0.645 0.610 78.73 67.25 73.80 73.26

### v2

@@ Abstract
Proprietary LMs such as GPT-4 are often employed to assess the quality of responses from various LMs. However, concerns including transparency, controllability, and affordability strongly motivate the development of open-source LMs specialized in evaluations. On the other hand, existing open evaluator LMs exhibit critical shortcomings: 1) they issue scores that significantly diverge from those assigned by humans, and 2) they lack the flexibility to perform both direct assessment and pairwise ranking, the two most prevalent forms of assessment. Additionally, they often do not possess the ability to evaluate based on custom evaluation criteria, focusing instead on general attributes like helpfulness and harmlessness. To address these issues, we introduce Prometheus 2. Prometheus 2 is more powerful than its predecessor, and closely mirrors human and GPT-4 judgements. Moreover, it is capable of processing both direct assessment and pair-wise ranking formats grouped with a user-defined evaluation criteria. On four direct assessment benchmarks and four pairwise ranking benchmarks, Prometheus 2 scores the highest correlation and agreement with humans and proprietary LM judges among all tested open evaluator LMs. Our models, code, and data are all publicly available.  [^1] https://github.com/prometheus-eval/prometheus-eval
@@ 1 Introduction
However, relying on proprietary LMs for evaluation poses significant challenges. The lack of transparency about their training data compromises both fairness and reproducibility, making it problematic to use them in evaluation pipelines. Additionally, concerns regarding controllability and affordability also persist (Kim et al., 2023). To address these issues, recent works have focused on developing evaluator LMs that are open-access, transparent, and controllable (Kim et al., 2023; Wang et al., 2023a, b; Li et al., 2023a; Zhu et al., 2023; Jiang et al., 2023b, c; Lee et al., 2024). Yet, these models often yield scoring decisions that do not correlate well enough with human judgments or those made by proprietary LMs, failing to effectively simulate them. Moreover, open evaluator LMs are not flexible since they are typically trained only to perform either direct assessment or pairwise ranking and assess based on general public preferences like helpfulness and harmlessness, limiting their ability to handle diverse real-life scenarios.
@@ 2.1 Language Model-based Evaluation
To assess the generation capabilities of LMs, prior works such as the GEM benchmark (Gehrmann et al., 2021, 2022) employed ROUGE (Lin, 2004), BLEU (Papineni et al., 2002), and BERTScore (Zhang et al., 2019) as their metrics, which measure the lexical or semantic similarity between a reference answer and a response. However, these conventional metrics are prone to false negatives because they are not expressive enough to recognize responses that are of good quality but differ from the reference answer (Schluter, 2017; Freitag et al., 2020; Hanna and Bojar, 2021).
Recently, employing language models as a judge has gained attention as a promising paradigm to mimic the depth and granularity that human evaluation offers (Zheng et al., 2023; Liu et al., 2023b; Li et al., 2023b; Chan et al., 2023; Ye et al., 2023). To reduce the over-reliance on proprietary LMs, follow-up works suggest training language models specialized in evaluations (Cui et al., 2023; Kim et al., 2023; Jiang et al., 2023b, c; Li et al., 2023a; Lee et al., 2024). Yet, open evaluator LMs do not possess the flexibility to function in different evaluation schemes and show weak evaluation performance compared to proprietary LMs. We aim to bridge this gap by introducing Prometheus 2.
@@ 2.2 Weight Merging
Prior works have demonstrated that weight merging can enhance performance across various domains, including language modeling (Li et al., 2022; Matena and Raffel, 2022; Ilharco et al., 2022; Don-Yehiya et al., 2022; Gururangan et al., 2023; Yadav et al., 2024; Sukhbaatar et al., 2024), instruction-tuning (Jang et al., 2023b; Yu et al., 2023), and aligning to user preferences (Jang et al., 2023a; Rame et al., 2024; Wang et al., 2024). In our work, we specifically focus on enhancing the evaluation capabilities of open evaluator LMs. By merging models trained on different assessment formats—specifically, direct assessment and pairwise ranking—we aim to obtain an evaluator LM that not only functions in both formats but also shows as good evaluation performances as proprietary LMs.
@@ 3 Methodology
We propose a new recipe for training a unified evaluator LM based on merging the weights of models trained for direct assessment and pairwise ranking. We begin with background on direct assessment and pairwise ranking for evaluator LMs (Section 3.1, 3.2), followed by the construction process of our training data (Section 3.3). Finally, we present our methods to train state-of-the-art evaluator LMs, Prometheus 2 models (Section 3.4).
@@ 3.1 Direct Assessment
Direct assessment is mapping an instruction $i$ and response $r$ into a scalar value score $s$, such as ${f}_{direct}:(i,r)\mapsto s\text{ where }s\in\mathbb{R}$. For the scoring range, we use an integer between 1 and 5.
Prior works have identified several recipes to align the scores provided by evaluator LMs (${s}_{LM}$) and the scores assigned by humans (${s}_{human}$). For instance, Liu et al. (2023a) and Zheng et al. (2023) have shown that it is crucial to add a reference answer $a$ as input to the evaluator LM to maximize the correlation between ${s}_{LM}$ and ${s}_{human}$. Also, Zheng et al. (2023) and Ye et al. (2023) showed that prompting the language model to write verbal feedback ${v}_{r}$ before $s$ also improves the correlation between ${s}_{LM}$ and ${s}_{human}$. Lastly, Ye et al. (2023) and Kim et al. (2023) showed that by explicitly integrating evaluation criteria $e$, users can define the standards for model assessment, ensuring evaluations are flexible to specific needs rather than generic qualities. Specifically, $e$ is represented as a score rubric including a description for the criterion itself and a set of descriptions for each score between the scoring range. This is expressed as:
@@ 3.2 Pairwise Ranking
In pairwise ranking, the evaluation criterion $e$ does not include a set of descriptions for each score; instead, only the description of the evaluation criterion itself. Also, it is noteworthy that the verbal feedback ${v}_{{r}_{m},{r}_{n}}$ compares the commonalities and differences between ${r}_{m}$ and ${r}_{n}$ concerning $e$.
@@ 3.3 The Preference Collection
Popular pairwise ranking datasets such as HH-RLHF (Bai et al., 2022) or Ultra Feedback (Cui et al., 2023) do not include an evaluation criterion $e$ and a verbal feedback ${v}_{{r}_{m},{r}_{n}}$. To train an evaluator LM that could assess based on such criteria, we construct the Preference Collection, including 1K evaluation criteria. We apply two modifications to the Feedback Collection. First, since the Feedback Collection includes five responses for each instruction, each corresponding to a scoring decision between 1 and 5, we pair two out of the five responses, resulting in a total of ten combinations per instruction. Using the existing scoring decisions for each response, we determine which response is better and assign a new scoring decision for that pair (i.e., “Response A is better” or “Response B is better”). Second, to generate new verbal feedback ${v}_{{r}_{m},{r}_{n}}$ for each pair of responses, we prompt GPT-4-1106 to identify the commonalities and differences between the two responses.
The statistics of the resulting dataset are listed in Table 1 along with the Feedback Collection. We explain about our quality verification process of the Preference Collection in Appendix A. Also, we include the prompts we use for the augmentation process in Appendix H.
Evaluation Method Benchmark Metrics Judgment Source Reference Answer # Score Rubrics # Instructions # Judgments
Direct Assessment Vicuna Bench Correlation Proprietary LMs Y 80 80 320
MT Bench Correlation Proprietary LMs Y 80 80 320
FLASK Correlation Proprietary LMs & Humans Y 12 200 2,000
Feedback Bench Correlation Proprietary LMs Y 200 200 1,000
Pairwise Ranking HHH Align. Accuracy Humans N 4 221 221
MT Bench Human Judg. Accuracy Humans N 1 80 3,360
Auto-J Eval Accuracy Humans N 1 58 1,392
Preference Bench Accuracy Proprietary LMs Y 200 200 2,000
@@
3.4 Training Methods & Baselines
@@
Prompting involves querying an LM to make judgments in a specified evaluation format without training. We employ Llama-2-Chat-7,13,70B (Touvron et al., 2023); Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a); and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024) as our baselines. It’s worth noting that models not explicitly trained on feedback data often fail to generate responses in the required format, making it extremely difficult to parse scoring decisions. Although it is impractical for regular use, we make a fair comparison by infinitely looping until scores can be parsed. Also, we include proprietary LMs such as GPT-3.5-Turbo-0613; GPT-4-1106; and Claude-3-Opus.
@@
Single-Format training involves training a base model $\theta$ on either on a direct assessment feedback dataset ${D}_{d}$ or a pairwise ranking feedback dataset ${D}_{p}$. For single-format trained evaluator LMs, we test Prometheus-7,13B (Kim et al., 2023) (direct assessment); UltraRM-13B (Cui et al., 2023) (pairwise ranking); and PairRM-0.4B (Jiang et al., 2023c) (pairwise ranking). In addition, we also report the performances of single-format training Mistral-7B-Instruct-v0.2 and Mixtral-8x7B-Instruct-v0.1 on either direct assessment or pairwise ranking.
@@
Joint training involves training a base model $\theta$ on both a direct assessment feedback dataset ${D}_{d}$ and a pairwise ranking feedback dataset ${D}_{p}$. This enables the resulting evaluator LM to function across both evaluation formats. For jointly trained evaluator LMs, we test Auto-J (Li et al., 2023a). In addition, we report the performances of jointly training Mistral-7B and Mixtral-8x7B on both direct assessment and pairwise ranking.
@@
Weight Merging involves training two models, ${\theta}_{d}$ and ${\theta}_{p}$, separately on a direct assessment feedback dataset ${D}_{d}$ and a pairwise ranking feedback dataset ${D}_{p}$. Then, the final evaluator LM ${\theta}_{final}$ is obtained by merging ${\theta}_{d}$ and ${\theta}_{p}$. For example, linear merging is as follows:
@@
In addition to linear merging, we test 5 additional variants, namely Task Arithmetic merging (Ilharco et al., 2022), TIES merging (Yadav et al., 2024), DARE-TIES and DARE-Linear merging (Yu et al., 2023), and SLERP merging (Goddard et al., 2024). We include an explanation of these merging methods and ablation experiment results of the performance differences in Appendix G. Among them, DARE-Linear showed the best performance, and hence we used it to train the Prometheus 2 (7B & 8x7B) models. Details on the hyper-parameters for training and inference along with the prompt templates are all listed in Appendix B, I, J.
@@ 4 Experimental Setup
The statistics of all the benchmarks are in Table 2. The four direct assessment benchmarks are:
@@
In direct assessment, we conduct reference-based evaluations by appending the reference answer as the input. We use Pearson, Spearman, and Kendall-Tau as performance metrics to measure scoring correlations against reference evaluators. Moreover, we include the results of the reference-free direct assessment evaluation in Appendix F.
@@ 5 Experimental Results
In this section, we compare the evaluation capabilities of Prometheus-2 models with other baselines using a direct assessment format (Section 5.1) and a pairwise ranking format (Section 5.2). Additionally, we measure the consistency of the scores from evaluator LMs in Appendix E.
@@ 5.1 Direct Assessment Results
The direct assessment results are shown in Table 3. The scoring decisions of Prometheus 2 models (7B & 8x7B), GPT-4-1106, Claude-3-Opus, and human evaluators all strongly correlate with each other, yielding Pearson correlations higher than 0.5 regardless of the reference evaluator and benchmark. On the other hand, base LMs, single-format trained LMs, and jointly trained LMs show lower correlations, mostly falling below 0.5.
Notably, Prometheus 2 models outperform Prometheus and Auto-J by at least 0.2 units across benchmarks in their correlation with proprietary LMs. Moreover, on the FLASK benchmark, while the correlation between humans and GPT-4 is 0.679, the highest correlation previously achieved by Prometheus-13B with humans was 0.449. Prometheus-2-8x7B achieves a correlation of 0.555 with humans, halving the gap.
@@ 5.2 Pairwise Ranking Results
The pairwise ranking results are shown in Table 4. We exclude the results of Pair RM and Ultra RM on ‘w/ Tie’ settings since they could not process it.
@@
Training Data Evaluation Format Direct Assessment Benchmarks Pairwise Ranking Benchmarks
@@
Direct Assessment & Pairwise Ranking 0.666 0.548 0.659 0.624 74.66 70.78 75.07 73.50